# Jet-like structures in low-mass binary neutron star merger remnants

### v1

@@ Abstract
GW170817 and GRB 170817A provided direct evidence that binary neutron star (NSNS) mergers can produce short gamma-ray bursts (sGRBs). However, questions remain about the nature of the central engine. Depending on the mass, the remnant from a NSNS merger may promptly collapse to a black hole (BH), form a hypermassive neutron star (HMNS) which undergoes a delayed collapse to a BH, a supramassive neutron star (SMNS) with a much longer lifetime, or an indefinitely stable NS with a mass below the TOV limit. There is strong evidence that a BH with an accretion disk can launch a sGRB-compatible jet via the Blandford–Znajek mechanism, but whether a supramassive star can do the same is less clear. We have performed general relativistic magnetohydrodynamics (GRMHD) simulations of the merger of both irrotational and spinning, equal-mass NSNSs constructed from a piecewise polytropic representation of the nuclear SLy equation of state, with a range of gravitational (ADM) masses that yield remnants with mass above and below the supramassive limit. Each NS is endowed with a dipolar magnetic field extending from the interior into the exterior, as in a radio pulsar. We examine cases with different initial binary masses, including a case which produces a HMNS which collapses to a BH, and lower mass binaries that produce SMNS remnants. We find similar jet-like structures (helical magnetic field structures, a magnetically dominated evacuated funnel and mildly relativistic outflow from the poles) for both the SMNS and HMNS remnants that meet our basic critera for an incipient jet. The outflow for the HMNS case is consistent with a Blandford-Znajek (BZ) jet. There is sufficient evidence that such BZ-powered outflows can break out and produce ulrarelativistic jets so that we can describe the HMNS system as a sGRB progenitor. However, the incipient jets from the SMNS remnants have much more baryon pollution and we see indications of inefficient outflow acceleration and mixing with the surrounding debris torus. Therefore, we cannot conclude that outflows from SMNSs are the progenitors of sGRBs.
@@ I Introduction
The gravitational wave (GW) event GW170817 Abbott et al. (2017a) observed by the LIGO-Virgo consortium was remarkable both as the first GW signal identified as a NSNS merger, and because of the coincident observation of a short $\gamma$-ray burst (sGRB) GRB 170817A Abbott et al. (2017b); Coulter et al. (2017); Savchenko et al. (2017); Lazzati et al. (2017); Murguia-Berthier et al. (2017); Duffell et al. (2018); Lamb et al. (2018, 2019); Mooley et al. (2018a, b); Wu and MacFadyen (2018) followed by observations accross the electromagnetic spectrum. This ushered in the era of multimessenger (GW + electromagnetic (EM)) astronomy and provided direct evidence that sGRBs (or at least a subset of them) come from compact binary mergers where at least one of the companions is a neutron star (NS), as predicted by Paczynski (1986); Goodman (1986); Eichler et al. (1989); Narayan et al. (1992); Fox et al. (2005); D’Avanzo (2015). The total mass of the progenitor system of GW170817 is constrained to $(2.73,3.29)M_{\odot}$ with 90% confidence. The individual masses $m_{1},m_{2}$ of the binary components were inferred as $m_{1}\in(1.36,2.26)M_{\odot}$ and $m_{2}\in(0.86,1.36)M_{\odot}$, with the uncertainty being due to the degeneracy between the mass ratio and the aligned spin components Abbott et al. (2017a). These values are within the observed range of NS masses Kiziltan et al. (2013); Landry and Read (2021), while below the lower limit of the BH mass distribution as inferred from X-ray binary observations and formation from stellar collapse Ozel et al. (2010); Kreidberg et al. (2012). While mechanisms have been proposed for the formation of BHs with smaller masses (see Yang et al. (2018); Carr et al. (2021)), the observational evidence for BHs with masses below $3M_{\odot}$ is very weak, supporting the identification of GW170817 as a NSNS merger.
@@
The origin of the $\gamma$-ray emission in sGRBs is thought to be a narrowly collimated highly relativistic jet, 111where by “jet” we mean a collimated outflow of EM fields and plasma along the rotational axis of the source Young (1991). with Lorentz factors of at least $\Gamma\gtrsim 20$ Ghirlanda et al. (2018) with typical values being $\Gamma\sim O(10^{2})$ Granot and Kumar (2006); Zou and Piran (2010); Paczynski (1986); Ghirlanda et al. (2018). The burst is characterized by hard prompt emission arising from either internal shocks between shells within the jet Rees and Meszaros (1994); Piran (2004) or external shocks from the leading shell Katz (1994), at distances $\gtrsim 10^{5}$km Veres et al. (2018).
@@
It has been suggested that the additional soft thermal $\gamma$-ray component of the GRB 170817 emission originates from the hot dense cocoon surrounding the jet Goldstein et al. (2017). The unusually low luminosity of GRB 170817 ($L\sim 10^{47}$ erg s-1 compared to a typical $\gamma$-ray isotropic-equivalent luminosity of $\sim 10^{49}-10^{53}$ erg s-1 Li et al. (2016); Beniamini et al. (2020); Shapiro (2017)), the spectral lag of the afterglow, and radio emission consistent with superluminal apparent motion Fraija et al. (2019); Mooley et al. (2018a), have been attributed to its jet being viewed $\sim 20-30^{\circ}$ off-axis with a half-opening angle of the jet core of $\lesssim 5^{\circ}$ (see e.g Ghirlanda et al. (2019); Troja et al. (2020)). Indeed it has been suggested that the $\gamma$-ray flux we observed from GRB 170817 is from a sheath of slower material surrounding the jet core rather than the core itself due to the oblique viewing angle Ghirlanda et al. (2019), while the on-axis emission from the core likely has an isotropic-equivalent luminosity of $\geq 10^{52}$ erg s-1, similar to other sGRBs, powered by a kinetic energy outflow of $\gtrsim 10^{53}$ erg s-1. The late time non-thermal $X$-ray and radio afterglow has been attributed to the interaction with the interstellar medium and the production of synchrotron radiation at the external forward-shock Fraija et al. (2019), while the UV/optical/infrared transient is consistent with the kilonova or macronova model Cowperthwaite et al. (2017); Smartt et al. (2017); Kasliwal et al. (2017): a thermal mostly-isotropic transient powered by the radioactive decay of unstable nuclei formed from rapid neutron capture (the r-process) in the neutron-rich non-relativistic ejecta (see Metzger (2020) for a review). The ejecta mass has been estimated as $0.04\pm 0.01M_{\odot}$ ($\sim 1.4\%$ of the total binary mass) with velocities of $\sim 0.1c$ and $\sim 0.3c$ for the red and blue components respectively.
@@
While the estimated masses of the binary companions strongly suggest that GW170817 represents a NSNS merger, its post-merger fate is uncertain. In principle, a NSNS merger can produce one of four possible outcomes depending on the equation of state (EOS), the mass, and the spin of the post-merger remnant Shibata and Taniguchi (2006): i) If the mass is below the Tolman–Oppenheimer–Volkoff (TOV) limit, the maximum mass $M_{\textup{TOV}}$ for a zero-temperature nonrotating NS, then the merger remnant will live for a very long-time as a spinning NS. In the presence of a dissipative process, e.g., from pulsar magnetic dipole emission, angular momentum is removed, and the NS ultimately spins down (in a timescale of $\sim O(10^{3})$s for magnetic dipole radiation Shapiro and Teukolsky (1983); Thompson et al. (2004)) to a nonrotating stable NS. ii) Remnants with mass larger than $M_{\textup{TOV}}$ but smaller than the maximum mass for a uniformly rotating zero-temperature NS, $M_{\textup{sup}}$, are termed supramassive neutron stars (SMNSs) Cook et al. (1992) (see Fig. 1). Similar to (i), in the presence of a dissipative process, SMNSs also spin down, but this time the endpoint of their evolution is a BH instead of stable NS Friedman et al. (1988). iii) For masses larger than $M_{\textup{sup}}$ a metastable hypermassive neutron star (HMNS) forms that can be supported only by differential rotation Baumgarte et al. (2000). The hypermassive star persists for many orbital periods, typically $O(10)$ms, before collapsing to a BH (see e.g. Duez et al. (2004); Ruiz et al. (2018a)), after the differential rotation is lost through viscous effects, magnetic winding, and GWs. iv) Finally, for total initial binary masses above some dynamically determined threshold, $M_{\textup{thresh}}$, which depends on the EOS and the initial NS spin, the remnant undergoes prompt collapse to BH on a timescale of only a few ms Köppel et al. (2019); Kölsch et al. (2022).
@@
A key open question in ascertaining the fate of the merger in event GW170817, as well as the central engine behind sGRBs, is whether an ultarelativistic jet can only be powered by a BH with an accretion disk, or whether it can also be powered by a long-lived, highly magnetized NS remnant immersed into a gaseous environment of tidal debris Usov (1992); Duncan and Thompson (1992); Zhang and Meszaros (2001); Metzger et al. (2008, 2011); Bucciantini et al. (2012); Metzger and Piro (2014); Ruiz et al. (2018a); Ciolfi (2020). Accretion onto a spinning BH can power an ultrarelativistic jet through either the Blandford-Znajek (BZ) mechanism Blandford and Znajek (1977), $\nu-\bar{\nu}$ annihilation along the BH spin axis Eichler et al. (1989); Ruffert and Janka (1998), or a combination of both, although the BZ mechanism is more likely to produce the luminosity consistent with observed sGRBs Kyutoku et al. (2018). BHs that form from the prompt collapse of merger remnants are unlikely to produce jets, as there is not enough time for the magnetic field to grow to force-free values above the BH poles Ruiz and Shapiro (2017), a requirement of the BZ mechanism. On the other hand, the more highly magnetized accretion disk that forms around the BH after the collapse of the HMNS creates the optimal conditions for a BZ-driven jet Ruiz et al. (2016); Usov (1992).
@@
Jet formation from NS remnants however remains an open problem. These NS remnants generally do not have the required ergosphere for the BZ mechanism, [^2] However, see Tsokaros et al. (2019a, 2020a) for extreme examples of NSs with an ergosphere. however the large reservoir of rotational energy of the star ($\sim 10^{53}\textup{erg}$) can in principle be sufficient to power a sGRB, if it can be efficiently extracted via magnetic processes Usov (1992); Ciolfi (2018); Ciolfi et al. (2019); Bucciantini et al. (2012). Slowly decaying “X-ray plateaus”, lasting $10^{2}-10^{5}$s, in the soft X-ray afterglow of a subset of sGRBs have been cited as evidence for continuous energy injection from a magnetar central engine Ciolfi (2018); Rowlinson et al. (2013); Zhang and Meszaros (2001) (although no such plateau was observed in the afterglow of GRB 170817 Abbott et al. (2017b); Margalit and Metzger (2017)). The timescales are significantly larger than the accretion timescale for a stellar mass BH, while the X-ray emission could be explained by spin down radiation from a NS Jordana-Mitjans et al. (2022). However, several alternative explanations (e.g. Strang and Melatos (2019); Oganesyan et al. (2020); Dereli-Bégué et al. (2022); Ciolfi and Siegel (2015)) for these features have been proposed which are compatible with the BH + disk model. The main challenge for the magnetar central engine model is the requirement for a relatively baryon-free environment to launch the jet. Neutrino radiation from the NS remnant and / or magnetic processes produce an isotropic baryon wind in addition to the dynamical ejecta Dessart et al. (2009); Siegel and Metzger (2017), and the resulting baryon pollution may limit the maximum terminal Lorentz factor of the jet to $\mathcal{O}(10)$ Sarin and Lasky (2021), less than the typical $\gtrsim 100$ which is required. It has also been argued that if a jet does launch from a magnetar remnant it needs to do so within $\leq 100$ms post-merger to avoid the jet becoming choked by the wind Murguia-Berthier et al. (2014).
@@
Several different groups have conducted general relativistic magnetohydrodynamic (GRMHD) simulations of NSNS mergers over the last two decades (for instance Price and Rosswog (2006); Duez et al. (2006); Anderson et al. (2008); Giacomazzo et al. (2011); Rezzolla et al. (2011); Neilsen et al. (2014); Palenzuela et al. (2015); Kiuchi et al. (2015); Sekiguchi et al. (2015); Ruiz et al. (2016); Kawamura et al. (2016); Kiuchi et al. (2018); Dietrich et al. (2018); Ruiz et al. (2019); Ciolfi et al. (2019); Ruiz et al. (2020a); Ciolfi (2020); Mösta et al. (2020); Ruiz et al. (2021); Sun et al. (2022); Kiuchi et al. (2023); Zappa et al. (2023); Foucart et al. (2023)) each with different strengths and weaknesses. Simulations performed by our group (e.g. Ruiz et al. (2016)) demonstrated the formation of a collimated, mildly relativistic outflow, with a tightly wound helical magnetic field from the poles, powered by the BZ process from a BH with a magnetized accretion disk formed following the collapse of a HMNS remnant Ruiz et al. (2016). We identified this outflow as an incipient jet. The NSs were irrotational, equal mass and modelled with a polytropic $\Gamma=2$ EOS, where $\Gamma$ is the polytropic index. Similar results were obtained from simulations of black hole-neutron star (BHNS) mergers Paschalidis et al. (2015); Ruiz et al. (2018b) which also result in a BH surrounded by a magnetized accretion disk. Our studies were later followed up with simulations where the stars had initial spin Ruiz et al. (2019), different orientations of the magnetic dipole moment Ruiz et al. (2020a), realistic piecewise polytropic EOSs (SLy and H4) Ruiz et al. (2021), and simulations incorporating a M1 neutrino transport scheme Sun et al. (2022). These studies showed:
@@
In this work, we further this discussion by conducting a systematic investigation of GRMHD simulations of NSNS mergers with initial magnetic fields extending from the interior into the exterior of the NSs, with a range of gravitational (ADM) binary masses and different spins. By fixing the EOS, we investigate binary mergers that yield remnants with mass above and below the supramassive limit, resulting either in HMNS or SMNS rapidly rotating remnants. By employing both irrotational and spinning binaries we probe the effects of spin in incipient jet launching, expanding our previous studies Ruiz et al. (2019) to the supramassive regime. [^4] Note here that GRMHD studies of accretion disks around BHs (e.g.McKinney and Blandford (2009)) suggest that a rotating BH ($a/M\gtrsim 0.4$) is a necessary, but not sufficient, condition to produce a highly relativistic ($\Gamma\gtrsim 3$) jet. Therefore the spin of the merger remnant (either NS or BH) may be crucial to the existence or not of a jet.

### v2

@@ Abstract
GW170817 and GRB 170817A provided direct evidence that binary neutron star (NSNS) mergers can produce short gamma-ray bursts (sGRBs). However, questions remain about the nature of the central engine. Depending on the mass, the remnant from a NSNS merger may promptly collapse to a black hole (BH), form a hypermassive neutron star (HMNS) which undergoes a delayed collapse to a BH, a supramassive neutron star (SMNS) with a much longer lifetime, or an indefinitely stable NS with a mass below the TOV limit. There is strong evidence that a BH with an accretion disk can launch a sGRB-compatible jet via the Blandford-Znajek mechanism, but whether a supramassive star can do the same is less clear. We have performed general relativistic magnetohydrodynamics simulations of the merger of both irrotational and spinning, equal-mass NSNSs constructed from a piecewise polytropic representation of the nuclear SLy equation of state, with a range of gravitational masses that yield remnants with mass above and below the supramassive limit. Each NS is endowed with a dipolar magnetic field extending from the interior into the exterior, as in a radio pulsar. We examine cases with different initial binary masses, including a case which produces a HMNS which collapses to a BH, and lower mass binaries that produce SMNS remnants. We find similar jetlike structures (helical magnetic field structures, a magnetically dominated evacuated funnel, and mildly relativistic outflow from the poles) for both the SMNS and HMNS remnants that meet our basic criteria for an incipient jet. The outflow for the HMNS case is consistent with a Blandford-Znajek (BZ) jet. There is sufficient evidence that such BZ-powered outflows can break out and produce ulrarelativistic jets so that we can describe the HMNS system as a sGRB progenitor. However, the incipient jets from the SMNS remnants have much more baryon pollution and we see indications of inefficient outflow acceleration and mixing with the surrounding debris torus. Therefore, we cannot conclude that outflows from SMNSs are the progenitors of sGRBs.
@@ I Introduction
The gravitational wave (GW) event GW170817 Abbott et al. (2017a) observed by the LIGO-Virgo consortium was remarkable both as the first GW signal identified as a NSNS merger, and because of the coincident observation of a short $\gamma$-ray burst (sGRB) GRB 170817A Abbott et al. (2017b); Coulter et al. (2017); Savchenko et al. (2017); Lazzati et al. (2017); Murguia-Berthier et al. (2017); Duffell et al. (2018); Lamb et al. (2018, 2019); Mooley et al. (2018a, b); Wu and MacFadyen (2018) followed by observations accross the electromagnetic spectrum. This ushered in the era of multimessenger [GW + electromagnetic (EM)] astronomy and provided direct evidence that sGRBs (or at least a subset of them) come from compact binary mergers where at least one of the companions is a neutron star (NS), as predicted by Paczynski (1986); Goodman (1986); Eichler et al. (1989); Narayan et al. (1992); Fox et al. (2005); D’Avanzo (2015). The total mass of the progenitor system of GW170817 is constrained to $(2.73,3.29)M_{\odot}$ with 90% confidence. The individual masses $m_{1},m_{2}$ of the binary components were inferred as $m_{1}\in(1.36,2.26)M_{\odot}$ and $m_{2}\in(0.86,1.36)M_{\odot}$, with the uncertainty being due to the degeneracy between the mass ratio and the aligned spin components Abbott et al. (2017a). These values are within the observed range of NS masses Kiziltan et al. (2013); Landry and Read (2021), while below the lower limit of the BH mass distribution as inferred from X-ray binary observations and formation from stellar collapse Ozel et al. (2010); Kreidberg et al. (2012). While mechanisms have been proposed for the formation of BHs with smaller masses (see Yang et al. (2018); Carr et al. (2021)), the observational evidence for BHs with masses below $3M_{\odot}$ is very weak, supporting the identification of GW170817 as a NSNS merger.
@@
The origin of the $\gamma$-ray emission in sGRBs is thought to be a narrowly collimated highly relativistic jet, 111where by “jet” we mean a collimated outflow of EM fields and plasma along the rotational axis of the source Young (1991). with Lorentz factors of at least $\Gamma\gtrsim 20$ Ghirlanda et al. (2018) with typical values being $\Gamma\sim O(10^{2})$ Granot and Kumar (2006); Zou and Piran (2010); Paczynski (1986); Ghirlanda et al. (2018). The typical isotropic-equivalent $\gamma$-ray luminosity of observed sGRBs is $\sim 10^{49}-10^{54}$ erg s-1 Li et al. (2016); Beniamini et al. (2020); Shapiro (2017) corresponding to a real $\gamma$-ray luminosity of $10^{47}-10^{52}$ erg s-1 Zhu et al. (2023). The burst is characterized by hard prompt emission arising from either internal shocks between shells within the jet Rees and Meszaros (1994); Piran (2004) or external shocks from the leading shell Katz (1994), at distances $\gtrsim 10^{5}$ km Veres et al. (2018).
@@
It has been suggested that the additional soft thermal $\gamma$-ray component of the GRB 170817 emission originates from the hot dense cocoon surrounding the jet Goldstein et al. (2017). The unusually low isotropic-equivalent luminosity of GRB 170817 ($L_{\gamma,\textup{iso}}\sim 10^{47}$erg s-1), the spectral lag of the afterglow, and radio emission consistent with superluminal apparent motion Fraija et al. (2019); Mooley et al. (2018a), have been attributed to its jet being viewed $\sim 20-30^{\circ}$ off-axis with a half-opening angle of the jet core of $\lesssim 5^{\circ}$ (see e.g Ghirlanda et al. (2019); Troja et al. (2020)). Indeed it has been suggested that the $\gamma$-ray flux we observed from GRB 170817 is from a sheath of slower material surrounding the jet core rather than the core itself due to the oblique viewing angle Ghirlanda et al. (2019), while the on-axis emission from the core likely has an isotropic-equivalent luminosity of $\geq 10^{51}$ erg s-1, similar to other sGRBs, powered by a isotropic-equivalent kinetic energy outflow of $\gtrsim 10^{53}$ erg s-1. The late time non-thermal $X$-ray and radio afterglow has been attributed to the interaction with the interstellar medium and the production of synchrotron radiation at the external forward-shock Fraija et al. (2019), while the UV/optical/infrared transient is consistent with the kilonova or macronova model Cowperthwaite et al. (2017); Smartt et al. (2017); Kasliwal et al. (2017): a thermal mostly-isotropic transient powered by the radioactive decay of unstable nuclei formed from rapid neutron capture (the r-process) in the neutron-rich non-relativistic ejecta (see Metzger (2020) for a review). The ejecta mass has been estimated as $0.04\pm 0.01M_{\odot}$ ($\sim 1.4\%$ of the total binary mass) with velocities of $\sim 0.1c$ and $\sim 0.3c$ for the red and blue components respectively.
@@
While the estimated masses of the binary companions strongly suggest that GW170817 represents a NSNS merger, its post-merger fate is uncertain. In principle, a NSNS merger can produce one of four possible outcomes depending on the equation of state (EOS), the mass, and the spin of the post-merger remnant Shibata and Taniguchi (2006): i) If the mass is below the Tolman–Oppenheimer–Volkoff (TOV) limit, the maximum mass $M_{\textup{TOV}}$ for a zero-temperature nonrotating NS, then the merger remnant will live for a very long-time as a spinning NS. In the presence of a dissipative process, e.g. from pulsar magnetic dipole emission, angular momentum is removed, and the NS ultimately spins down to a nonrotating stable NS. The spindown timescale is $\sim 10^{3}B^{-2}_{15}\;T^{2}_{\textup{ms}}\;\textup{s}$ for magnetic dipole radiation Shapiro and Teukolsky (1983); Thompson et al. (2004); Lander and Jones (2018) where $B_{15}$ is the magnetic field strength in units of $10^{15}G$ and $T_{\textup{ms}}$ the rotation period in ms. ii) Remnants with mass larger than $M_{\textup{TOV}}$ but smaller than the maximum mass for a uniformly rotating zero-temperature NS, $M_{\textup{sup}}$, are termed supramassive neutron stars (SMNSs) Cook et al. (1992) (see Fig. 1). Similar to (i), in the presence of a dissipative process, SMNSs also spin down, but this time the endpoint of their evolution is a BH instead of stable NS Friedman et al. (1988). The SMNS lifetime depends similarly on $B$ and $T$, and is typically of the order $\sim 10^{3}$s Beniamini and Lu (2021), but its exact value depends on how close the star is to the turning point for uniformly rotating stars. iii) For masses larger than $M_{\textup{sup}}$ a metastable hypermassive neutron star (HMNS) forms that can be supported only by differential rotation Baumgarte et al. (2000). The hypermassive star persists for many orbital periods, typically $O(10)$ms, before collapsing to a BH (see e.g. Duez et al. (2004); Ruiz et al. (2018a)), after the differential rotation is lost through viscous effects, magnetic winding, and GWs. iv) Finally, for total initial binary masses above some dynamically determined threshold, $M_{\textup{thresh}}$, which depends on the EOS and the initial NS spin, the remnant undergoes prompt collapse to BH on a timescale of only a few ms Köppel et al. (2019); Kölsch et al. (2022).
@@
A key open question in ascertaining the fate of the merger in event GW170817, as well as the central engine behind sGRBs, is whether an ultarelativistic jet can only be powered by a BH with an accretion disk, or whether it can also be powered by a long-lived, highly magnetized NS remnant immersed into a gaseous environment of tidal debris Usov (1992); Duncan and Thompson (1992); Zhang and Meszaros (2001); Metzger et al. (2008, 2011); Bucciantini et al. (2012); Metzger and Piro (2014); Ruiz et al. (2018a); Ciolfi (2020). Accretion onto a spinning BH can power an ultrarelativistic jet through either the Blandford-Znajek (BZ) mechanism Blandford and Znajek (1977), neutrino-antineutrino annihilation along the BH spin axis Eichler et al. (1989); Ruffert and Janka (1998), or a combination of both, although the BZ mechanism is more likely to produce the luminosity consistent with observed sGRBs Kyutoku et al. (2018). BHs that form from the prompt collapse of merger remnants are unlikely to produce jets, as there is not enough time for the magnetic field to grow to force-free values above the BH poles Ruiz and Shapiro (2017), a requirement of the BZ mechanism. On the other hand, the more highly magnetized accretion disk that forms around the BH after the collapse of the HMNS creates the optimal conditions for a BZ-driven jet Ruiz et al. (2016); Usov (1992).
@@
Jet formation from NS remnants however remains an open problem. These NS remnants generally do not have the required ergosphere for the BZ mechanism, [^2] However, see Tsokaros et al. (2019a, 2020a) for extreme examples of NSs with an ergosphere. however the large reservoir of rotational energy of the star ($\sim 10^{53}\textup{erg}$) can in principle be sufficient to power a sGRB, if it can be efficiently extracted via magnetic processes Usov (1992); Ciolfi (2018); Ciolfi et al. (2019); Bucciantini et al. (2012). Slowly decaying “X-ray plateaus”, lasting $10^{2}-10^{5}$s, in the soft X-ray afterglow of a subset of sGRBs have been cited as evidence for continuous energy injection from a magnetar central engine Ciolfi (2018); Rowlinson et al. (2013); Zhang and Meszaros (2001) (although no such plateau was observed in the afterglow of GRB 170817 Abbott et al. (2017b); Margalit and Metzger (2017)). The timescales are significantly larger than the accretion timescale for a stellar mass BH, while the X-ray emission could be explained by spin down radiation from a NS Jordana-Mitjans et al. (2022). However, several alternative explanations (e.g. Strang and Melatos (2019); Oganesyan et al. (2020); Dereli-Bégué et al. (2022); Ciolfi and Siegel (2015)) for these features have been proposed which are compatible with the BH + disk model. The additional energy injection from a magnetar remnant increases the energy of the quasi-isotropic ejecta and the associated kilonova, producing stronger radio emission at late-times Metzger and Piro (2014). The non-detection of such late-time radio emission has been used to rule out a magnetar remnant central engine for some sGRBs Metzger and Piro (2014); Horesh et al. (2016); Fong et al. (2016). Conversely, other authors have argued that, given the uncertainties in the physical parameters, radio observations of sGRB afterglows remain broadly compatible with magnetar remnants Liu et al. (2019).
The main challenge for the magnetar central engine model is the requirement for a relatively baryon-free environment to launch the jet. Neutrino radiation from the NS remnant and / or magnetic processes produce an isotropic baryon wind in addition to the dynamical ejecta Dessart et al. (2009); Siegel and Metzger (2017), and the resulting baryon pollution may limit the maximum terminal Lorentz factor of the jet to $\mathcal{O}(10)$ Sarin and Lasky (2021), less than the typical $\gtrsim 100$ which is required. It has also been argued that if a jet does launch from a magnetar remnant it needs to do so within $\leq 100$ms post-merger to avoid the jet becoming choked by the wind Murguia-Berthier et al. (2014).
@@
Several different groups have conducted general relativistic magnetohydrodynamic (GRMHD) simulations of NSNS mergers over the last two decades (for instance Price and Rosswog (2006); Duez et al. (2006); Anderson et al. (2008); Giacomazzo et al. (2011); Rezzolla et al. (2011); Neilsen et al. (2014); Palenzuela et al. (2015); Kiuchi et al. (2015); Sekiguchi et al. (2015); Ruiz et al. (2016); Kawamura et al. (2016); Kiuchi et al. (2018); Dietrich et al. (2018); Ruiz et al. (2019); Ciolfi et al. (2019); Ruiz et al. (2020a); Ciolfi (2020); Mösta et al. (2020); Ruiz et al. (2021); Sun et al. (2022); Kiuchi et al. (2023); Zappa et al. (2023); Foucart et al. (2023); Most and Quataert (2023); Radice and Bernuzzi (2023); Combi and Siegel (2023); Kiuchi et al. (2024); Aguilera-Miret et al. (2023); Zenati et al. (2024)) each with different strengths and weaknesses. Simulations performed by our group (e.g. Ruiz et al. (2016)) demonstrated the formation of a collimated, mildly relativistic outflow, with a tightly wound helical magnetic field from the poles, powered by the BZ process from a BH with a magnetized accretion disk formed following the collapse of a HMNS remnant Ruiz et al. (2016). We identified this outflow as an incipient jet. The NSs were irrotational, equal mass and modeled with a polytropic $\Gamma=2$ EOS, where $\Gamma$ is the adiabatic index. Similar results were obtained from simulations of black hole-neutron star (BHNS) mergers Paschalidis et al. (2015); Ruiz et al. (2018b) which also result in a BH surrounded by a magnetized accretion disk. Our studies were later followed up with simulations where the stars had initial spin Ruiz et al. (2019), different orientations of the magnetic dipole moment Ruiz et al. (2020a), realistic piecewise polytropic EOSs (SLy and H4) Ruiz et al. (2021), and simulations incorporating a M1 neutrino transport scheme Sun et al. (2022). These studies showed:
@@
In this work, we further this discussion by conducting a systematic investigation of GRMHD simulations of NSNS mergers with initial magnetic fields extending from the interior into the exterior of the NSs, with a range of Arnowitt–Deser–Misner (ADM) gravitational binary masses and different spins. By fixing the EOS, we investigate binary mergers that yield remnants with mass above and below the supramassive limit, resulting either in HMNS or SMNS rapidly rotating remnants. By employing both irrotational and spinning binaries we probe the effects of spin in incipient jet launching, expanding our previous studies Ruiz et al. (2019) to the supramassive regime. [^4] Note here that GRMHD studies of accretion disks around BHs (e.g.McKinney and Blandford (2009)) suggest that a rotating BH ($a/M\gtrsim 0.4$) is a necessary, but not sufficient, condition to produce a highly relativistic ($\Gamma\gtrsim 3$) jet. Therefore the spin of the merger remnant (either NS or BH) may be crucial to the existence or not of a jet.